Lightfoot Inc., a software development firm, has stock outstanding as follows: 25,000 shares of cumulative preferred 3% stock, $25 par, and 31,000 shares of $50 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, $7,250; second year, $11,750; third year, $59,570; fourth year, $94,700.

Calculate the dividends per share on each class of stock for each of the four years. Round all answers to two decimal places. If no dividends are paid in a given year, enter "0".

1st Year 2nd Year 3rd Year 4th Year
Preferred stock (dividend per share) $ $ $ $
Common stock (dividend per share) $ $ $ $
Answer #1

Par value per preferred share

Dividend rate

Dividend per preferred share

Number of preferred shares

Preferred dividend

Annual preferred dividend

$25

3%

25 x 3% = $0.75

25,000

25,000 x 0.75 = 18,750

Total cash dividend paid

Paid to preferred

Paid to common

Dividend in arrears at year end

Year 1

7,250

7,250

0

11,500

Year 2

11,750

11,750

0

18,500

Year 3

59,570

37,250

22,320

0

Year 4

94,700

18,750

75,950

0

$173,270

$75,000

$98,270

0

Preferred stock (Dividend per share)

Common stock (Dividend per share)

First year

7,250/25,000 = $0.29

0

Second year

11,750/25,000 = $0.47

0

Third year

37,250/25,000 = $1.49

22,320/31,000 = $0.72

Fourth year

18,750/25,000 = $0.75

75,950/31,000 = $2.45

1st year

2nd year

3rd year

4th year

Preferred stock (Dividend per share)

$0.29

$0.47

$1.49

$0.75

Common stock (Dividend per share)

0

0

$0.72

$2.45

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.